# Large Multi-modality Model Assisted AI-Generated Image Quality Assessment

### v1

@@ Abstract.
Traditional deep neural network (DNN)-based image quality assessment (IQA) models leverage convolutional neural networks (CNN) or Transformer to learn the quality-aware feature representation, achieving commendable performance on natural scene images. However, when applied to AI-Generated images (AGIs), these DNN-based IQA models exhibit subpar performance. This situation is largely due to the semantic inaccuracies inherent in certain AGIs caused by uncontrollable nature of the generation process. Thus, the capability to discern semantic content becomes crucial for assessing the quality of AGIs. Traditional DNN-based IQA models, constrained by limited parameter complexity and training data, struggle to capture complex fine-grained semantic features, making it challenging to grasp the existence and coherence of semantic content of the entire image. To address the shortfall in semantic content perception of current IQA models, we introduce a large Multi-modality model Assisted AI-Generated Image Quality Assessment (MA-AGIQA) model, which utilizes semantically informed guidance to sense semantic information and extract semantic vectors through carefully designed text prompts. Moreover, it employs a mixture of experts (MoE) structure to dynamically integrate the semantic information with the quality-aware features extracted by traditional DNN-based IQA models. Comprehensive experiments conducted on two AI-generated content datasets, AIGCQA-20k and AGIQA-3k show that MA-AGIQA achieves state-of-the-art performance, and demonstrate its superior generalization capabilities on assessing the quality of AGIs. The code will be available.
@@ 1. Introduction
As shown in Figure 1, AI-generated images, derived from advanced image generative models such as generative adversarial networks (GANs) (Ma et al., 2023), diffision (Ho et al., 2020) and related variant (Rombach et al., 2022b; Xu et al., 2018; Ramesh et al., 2021, 2022; dreamlike art, 2023; DeepFloyd, 2023; Holz, 2023; PlaygroundAI, 2023), often exhibit issues not commonly found in naturally captured images. Visual quality of AGIs depends not only on basic visual features such as noise, blur (Su et al., 2021; Li et al., 2022; Zhang et al., 2024), etc., but also on more intricate semantic perception (Li et al., 2024a), such as existence of reasonable semantic content, scene plausibility, and the coherence among objects (Wu et al., 2023b; Zhang et al., 2023; Li et al., 2024b; Wu et al., 2023c, 2024b). Although re-training existing IQA models on AGIs datasets leads to improved outcomes, it fails to achieve optimal performance. One reason is that traditional DNN models, especially early convolutional neural networks (CNNs), despite their notable achievements in tasks like image recognition and classification (Simonyan and Zisserman, 2015; He et al., 2015; Szegedy et al., 2014), still struggle to grasp the fine-grained semantic content of images (Zhang et al., 2021). What’s more, traditional DNN-based IQA models fail to capture the intrinsic characteristics essential for assessing image quality and thus exhibit poor generalization abilities. Hence, we argue that the quality assessment models of AGIs are still in their infancy and need further exploration.
@@
In this paper, we introduce a large Multi-modality model Assisted AI-Generated Image Quality Assessment (MA-AGIQA) framework, which enhances the capacity of traditional DNN-based IQA models to understand semantic content by incorporating LMM. Our approach initially repurposes a DNN, MANIQA (Yang et al., 2022), as an extractor for quality-aware features and establishes it as the training backbone for the MA-AGIQA framework. Subsequently, we guide a LMM, mPLUG-Owl2 (Ye et al., 2023), to focus on fine-grained semantic information through meticulously crafted prompts. We then extract and store the last-layer hidden vector from mPLUG-Owl2, merging it with features extracted by MANIQA to infuse the model with rich semantic insights. Finally, we employ a MoE to dynamically integrate quality-aware features with fine-grained semantic features, catering to the unique focal points of different images. As demonstrated in Figure 2, our approach surpasses MANIQA in terms of SRCC, particularly within subsets comprising semantically rich images overflowing with graininess, indicating that our methodology shows remarkable congruence with the human visual system’s (HVS) perceptual capabilities. MA-AGIQA achieves SRCC values of 0.8939 and 0.8644 on the AGIQA-3k and AIGCQA-20k datasets, respectively, exceeding the state-of-the-art models by 2.03% and 1.37%, and also demonstrates superior cross-dataset performance.
@@
We evaluate the MA-AGIQA model on two AI-generated IQA datasets. Experimental results demonstrate that our model surpasses current state-of-the-art methods without extra training data and also showcases superior cross-dataset performance. Extensive ablation studies further validate the effectiveness of each component.
@@ 2. Related Work
With the advent of deep learning, CNNs have revolutionized many tasks in computer vision. (Kang et al., 2014) is pioneer in applying deep convolutional neural networks to NR-IQA. Its methodology employs CNNs to directly learn representations of image quality from raw image patches, bypassing the need for handcrafted features or a reference image. Following this, DBCNN (Zhang et al., 2018) introduces a deep bilinear CNN for blind image quality assessment (BIQA) (Zhai and Min, 2020), innovatively merging two CNN streams to address both synthetic and authentic image distortions separately. Furthermore, HyperIQA (Su et al., 2020), a self-adaptive hyper network, evaluates the quality of authentically distorted images through a novel three-stage process: content understanding, perception rule learning, and quality prediction.
The success of Vision Transformers (ViT) (Dosovitskiy et al., 2020) in various computer vision tasks has led to significant advancements. In the realm of IQA, IQT (You and Korhonen, 2021) leverages the combination of reference and distorted image features, extracted by CNNs, as inputs for a Transformer-based quality prediction task. MUSIQ (Ke et al., 2021) utilizes a Transformer to encode distortion image features across three scales, addressing the challenge of varying input image sizes during training and testing. TReS introduces relative ranking and self-consistency loss to capitalize on the abundant self-supervisory information available, aiming to decrease the network’s sensitivity. What’s more, MANIQA (Yang et al., 2022) explored multi-dimensional feature interaction, utilizing spatial and channel structural information to calculate a non-local representation of the image, enhancing the model’s ability to assess image quality comprehensively.
@@ 3. Method
As depicted in Figure 3, framework of MA-AGIQA is structured into three sections. Section 3.1 introduces our adoption of a DNN, specifically MANIQA (Yang et al., 2022), tailored for the AGIs quality assessment task, serving as our primary training backbone. In Section 3.2, we incorporate the LMM mPLUG-Owl2 (Ye et al., 2023) as a feature extractor. This component is crucial for acquiring fine-grained semantic features via carefully crafted text prompts. Lastly, Section 3.3 addresses the variability in focal points across different images. To adaptively integrate the feature vectors during training, we utilizes a MoE structure for feature fusion. This approach ensures that the most salient features are emphasized. Further details are elaborated below.
@@ 3.1. Quality-aware Feature Extraction
To leverage the capability of DNNs to adapt to the data distribution of specific tasks, we employ MANIQA (Yang et al., 2022) as a quality-aware feature extractor. MANIQA enhances the evaluation of image quality by applying attention mechanisms across both the channel and spatial dimensions, thereby increasing the interaction among various regions of the image, both globally and locally. This approach generates projections and () for a given image, and the final rating of the whole image is determined through the sum of multiplication of by , which can be illustrated as Equation (1):
@@
During the training phase, the parameters of modified MANIQA are continuously updated. This refinement process ensures that MANIQA can extract features more relevant to the quality of AGIs. Furthermore, the training process facilitates a more seamless integration between MANIQA and LMM, leading to superior outcomes.
@@ 3.2. Fine-grained Semantic Feature Extraction
LMMs are capable of understanding and analyzing the semantic content of images and their relationship with human cognition. They assess whether different parts of an image form a cohesive whole and evaluate whether the elements within the picture are semantically coherent (Gao et al., 2023; OpenAI, 2023; Liu et al., 2023). mPLUG-Owl2 (Ye et al., 2023) employs a modality-adaptive language decoder to handle different modalities within distinct modules, which mitigates the issue of modality interference. Given the importance of effectively guiding the model through textual prompts to elicit the desired output, we have selected mPLUG-Owl2 as our feature extractor.
@@
Consequently, we try to propose the rational design of prompts leading LMMs to obtain those image semantic content. mPLUG-Owl2 possess the ability to understand fine-grained semantic contents, but without carefully designed input prompts, some prompts, such as ”Please evaluate if the image quality is compromised due to violations of common human sense or logic?” although it expresses the desire for the model to assess whether the semantic content of the image contradicts human perception, would lead to unsatisfactory results. To better utilize mPLUG-Owl2 for the task of evaluating AGIs, we meticulously designed prompts to guide the LMM. Specifically, we designed two prompts, denoted as and respctively,
@@
corresponding to the existence of semantic content and the coherence of semantic content in images, respectively. Test results, as shown in Figure 5 using the mPLUG-Owl2 official demo111https://modelscope.cn/studios/iic/mPLUG-Owl2/summary, have proven these questions to be effective.
However, the textual output from mPLUG-Owl2 is not immediately conducive to being utilized by MANIQA to impart semantic insights. To bridge this gap, it’s essential to obtain the information provided by mPLUG-Owl2 into a format that MANIQA can easily leverage. SO we extract features from the final layer of mPLUG-Owl2’s hidden layers, achieving an accessible embedded representation of the LMM’s output. This output is a tensor with dimensions of [token_length, hidden_size], where ”token_length” represents the number of output tokens, and ”hidden_size” denotes the dimensionality of the hidden layer representations associated with each token. For mPLUG-Owl2, the hidden_size is set to 4096. Subsequently, we conduct an averaging operation across the token dimension, yielding a vector with dimensions 1x4096. This vector then serves as the basis for further feature fusion procedures. The process can be represented as Equation (3) :
@@ 3.3. Adaptive Fusion Module
Given the complex influence of color, composition, details, semantic content, and other factors on image quality, simply concatenating the extracted features may not always yield the best results. To dynamically fuse a variety of complementary features, we propose the adaptive fusion module (AFM) for organic feature integration. This process can be divided into two main parts. The first part involves transforming the extracted features into a unified vector space of the same dimension, allowing for vector fusion operations. Specifically, for features extracted by MANIQA, this transformation block applies a fully connected (Fc) layer, transforming them to the same dimension as the original features (1x784) to provide a richer combination. For features derived from mPLUG-Owl2, it uses a Fc layer to project them onto a 1x784 dimension, followed by a relu activation layer and a dropout layer to enhance the network’s expressive power and generalization. The second part employs a MoE to dynamically fuse the three features. The MoE’s gating network takes the transformed three features as input and outputs dynamic weights , corresponding to the three features’ contributions to image quality. Structurally, this gating network comprises a Fc layer and a sigmoid layer. The final image quality representation vector can be obtained through a weighted sum of the three feature vectors. Following the denotation which sign the three features as , , , this process can be represented as:
@@
| Type | Method | AGIQA-3k | AIGCQA-20K-Image | ||||||
| SRCC | PLCC | KRCC | RMSE | SRCC | PLCC | KRCC | RMSE | ||
| Handcrafted feature-based | BRISQUE (Mittal et al., 2012a) | 0.4726 | 0.5612 | 0.3227 | 0.8299 | 0.1663 | 0.3580 | 0.1112 | 0.6813 |
| NIQE (Mittal et al., 2012b) | 0.5236 | 0.5668 | 0.3637 | 0.8260 | 0.2085 | 0.3378 | 0.1394 | 0.6868 | |
| ILNIQE (Zhang et al., 2015) | 0.6097 | 0.6551 | 0.4318 | 0.7576 | 0.3359 | 0.4551 | 0.2290 | 0.6497 | |
| LMM-based | CLIPIQA (Wang et al., 2023) | 0.6524 | 0.6968 | 0.4632 | 0.7191 | 0.4147 | 0.6459 | 0.2861 | 0.5570 |
| CLIPIQA+ (Wang et al., 2023) | 0.6933 | 0.7493 | 0.4957 | 0.664 | 0.4553 | 0.6682 | 0.3169 | 0.5428 | |
| Q-Align (Wu et al., 2023a) | 0.6728 | 0.6910 | 0.4728 | 0.7204 | 0.6743 | 0.6815 | 0.4808 | 0.5199 | |
| Traditional DNN-based | HyperIQA (Su et al., 2020) | 0.8509 | 0.9049 | 0.6685 | 0.4134 | 0.8162 | 0.8329 | 0.6207 | 0.3902 |
| MANIQA (Yang et al., 2022) | 0.8618 | 0.9115 | 0.6839 | 0.4111 | 0.8507 | 0.8870 | 0.6612 | 0.3273 | |
| DBCNN (Zhang et al., 2018) | 0.8263 | 0.8900 | 0.6393 | 0.4533 | 0.8054 | 0.8483 | 0.6121 | 0.3726 | |
| StairIQA (Sun et al., 2023) | 0.8343 | 0.8933 | 0.6485 | 0.4510 | 0.7899 | 0.8428 | 0.6053 | 0.3927 | |
| BAID (Yi et al., 2023) | 0.1304 | 0.2030 | 0.0854 | 0.9487 | 0.1652 | 0.1483 | 0.1279 | 0.7297 | |
| MUSIQ (Ke et al., 2021) | 0.8261 | 0.8657 | 0.6400 | 0.4907 | 0.8329 | 0.8646 | 0.6403 | 0.3634 | |
| DL with LMM | MA-AGIQA | 0.8939 | 0.9273 | 0.7211 | 0.3756 | 0.8644 | 0.9050 | 0.6804 | 0.3104 |

### v2

@@ Abstract.
11footnotetext: Corresponding author.††footnotetext: All of the authors are with Department of Electronic Engineering, Shanghai Jiao Tong University, Shanghai 200240 China. Puyi Wang is also with the Zhiyuan College, Shanghai Jiao Tong University.Traditional deep neural network (DNN)-based image quality assessment (IQA) models leverage convolutional neural networks (CNN) or Transformer to learn the quality-aware feature representation, achieving commendable performance on natural scene images. However, when applied to AI-Generated images (AGIs), these DNN-based IQA models exhibit subpar performance. This situation is largely due to the semantic inaccuracies inherent in certain AGIs caused by uncontrollable nature of the generation process. Thus, the capability to discern semantic content becomes crucial for assessing the quality of AGIs. Traditional DNN-based IQA models, constrained by limited parameter complexity and training data, struggle to capture complex fine-grained semantic features, making it challenging to grasp the existence and coherence of semantic content of the entire image. To address the shortfall in semantic content perception of current IQA models, we introduce a large Multi-modality model Assisted AI-Generated Image Quality Assessment (MA-AGIQA) model, which utilizes semantically informed guidance to sense semantic information and extract semantic vectors through carefully designed text prompts. Moreover, it employs a mixture of experts (MoE) structure to dynamically integrate the semantic information with the quality-aware features extracted by traditional DNN-based IQA models. Comprehensive experiments conducted on two AI-generated content datasets and two traditional IQA datasets show that MA-AGIQA achieves state-of-the-art performance, and demonstrate its superior generalization capabilities on assessing the quality of AGIs. The code is available at https://github.com/wangpuyi/MA-AGIQA.
@@ 1. Introduction
As shown in Figure 1, AI-generated images, derived from advanced image generative models such as generative adversarial networks (GANs) (Ma et al., 2023), diffision (Ho et al., 2020) and related variant (Rombach et al., 2022b; Xu et al., 2018; Ramesh et al., 2021, 2022; dreamlike art, 2023; DeepFloyd, 2023; Holz, 2023; PlaygroundAI, 2023), often exhibit issues not commonly found in naturally captured images. Visual quality of AGIs depends not only on basic visual features such as noise, blur (Su et al., 2021; Li et al., 2022; Zhang et al., 2024), etc., but also on more intricate semantic perception (Li et al., 2024a), such as existence of reasonable semantic content, scene plausibility, and the coherence among objects (Wu et al., 2023b; Zhang et al., 2023; Li et al., 2024b; Wu et al., 2023c, 2024b). Although re-training existing IQA models on AGIs datasets leads to improved outcomes, it fails to achieve optimal performance. One reason is that traditional DNN models, especially early convolutional neural networks (CNNs), despite their notable achievements in tasks like image recognition and classification (Simonyan and Zisserman, 2015; He et al., 2015; Szegedy et al., 2014), still struggle to grasp the fine-grained semantic content of images (Zhang et al., 2021). What’s more, traditional DNN based IQA models fail to capture the intrinsic characteristics essential for assessing image quality and thus exhibit poor generalization abilities. Hence, we argue that the quality assessment models of AGIs are still in their infancy and need further exploration.
@@
In this paper, we introduce a large Multi-modality model Assisted AI-Generated Image Quality Assessment (MA-AGIQA) framework, which enhances the capacity of traditional DNN-based IQA models to understand semantic content by incorporating LMM. Our approach initially repurposes a DNN, MANIQA (Yang et al., 2022), as an extractor for quality-aware features and establishes it as the training backbone for the MA-AGIQA framework. Subsequently, we guide a LMM, mPLUG-Owl2 (Ye et al., 2023), to focus on fine-grained semantic information through meticulously crafted prompts. We then extract and store the last-layer hidden vector from mPLUG-Owl2, merging it with features extracted by MANIQA to infuse the model with rich semantic insights. Finally, we employ a MoE to dynamically integrate quality-aware features with fine-grained semantic features, catering to the unique focal points of different images.
As an evidence that our MA-AGIQA model are clearly superior to traditional DNN-based methods when it comes to AGIs, a subset with grainy images (the concept of ”grainy” is depicted like the grainy orange cat in Figure 1) is sampled out from test set of AIGCQA-20k. In this subset, MANIQA achieves an SRCC of 0.2545, which is 70.0% lower than its overall SRCC performance (0.8507). In contrast, our model achieves an SRCC of 0.8364 in this subset and has an overall performance of 0.8644. It demonstrates that our model possesses a significantly enhanced understanding of AGIs, particularly those whose quality is deeply intertwined with semantic elements. Besides, MA-AGIQA achieves SRCC values of 0.8939 and 0.8644 on the AGIQA-3k and AIGCQA-20k datasets, respectively, exceeding the state-of-the-art models by 2.03% and 1.37%, and also demonstrates superior cross-dataset performance.
@@
We evaluate the MA-AGIQA model on two AI-generated IQA datasets and two traditional IQA datasets. Experimental results demonstrate that our model surpasses current state-of-the-art methods without extra training data and also showcases superior cross-dataset performance. Extensive ablation studies further validate the effectiveness of each component.
@@ 2. Related Work
With the advent of deep learning, CNNs have revolutionized many tasks in computer vision. (Kang et al., 2014) is pioneer in applying deep convolutional neural networks to NR-IQA. Its methodology employs CNNs to directly learn representations of image quality from raw image patches, bypassing the need for handcrafted features or a reference image. Following this, DBCNN (Zhang et al., 2018) introduces a deep bilinear CNN for NR-IQA (Zhai and Min, 2020), innovatively merging two CNN streams to address both synthetic and authentic image distortions separately. Furthermore, HyperIQA (Su et al., 2020), a self-adaptive hyper network, evaluates the quality of authentically distorted images through a novel three-stage process: content understanding, perception rule learning, and quality prediction.
The success of Vision Transformers (ViT) (Dosovitskiy et al., 2020) in various computer vision tasks has led to significant advancements. In the realm of IQA, IQT (You and Korhonen, 2021) leverages the combination of reference and distorted image features, extracted by CNNs, as inputs for a Transformer-based quality prediction task. MUSIQ (Ke et al., 2021) utilizes a Transformer to encode distortion image features across three scales, addressing the challenge of varying input image sizes during training and testing. TReS (Golestaneh et al., 2022) introduces relative ranking and self-consistency loss to capitalize on the abundant self-supervisory information available, aiming to decrease the network’s sensitivity. What’s more, MANIQA (Yang et al., 2022) explored multi-dimensional feature interaction, utilizing spatial and channel structural information to calculate a non-local representation of the image, enhancing the model’s ability to assess image quality comprehensively.
@@ 3. Method
As depicted in Figure 2, framework of MA-AGIQA is structured into three sections. Section 3.1 introduces our adoption of a DNN, specifically MANIQA (Yang et al., 2022), tailored for the AGIs quality assessment task, serving as our primary training backbone. In Section 3.2, we incorporate the LMM mPLUG-Owl2 (Ye et al., 2023) as a feature extractor. This component is crucial for acquiring fine-grained semantic features via carefully crafted text prompts. Lastly, Section 3.3 utilizes a MoE structure for adaptive feature fusion, which addresses the variability in focal points across differenct images and ensures that the most salient features are emphasized. Further details are elaborated below.
@@ 3.1. Quality-aware Feature Extraction
To leverage the capability of DNNs to adapt to the data distribution of specific tasks, we employ MANIQA (Yang et al., 2022) as a quality-aware feature extractor. MANIQA enhances the image quality evaluation by applying attention mechanisms across both the channel and spatial dimensions, thereby increasing the interaction among various regions of the image, both globally and locally. This approach generates projections and () for a given image, and the final rating of the whole image is determined through the sum of multiplication of by , which can be illustrated as Equation (1):
@@
During the training phase, the parameters of modified MANIQA are continuously updated. This refinement process ensures that MANIQA can extract features more relevant to the quality of AGIs and facilitates a more seamless integration between MANIQA and LMM, leading to superior outcomes.
@@ 3.2. Fine-grained Semantic Feature Extraction
LMMs are utilized to assess whether different parts of an image form a cohesive whole and whether the elements within the picture are semantically coherent (Gao et al., 2023; OpenAI, 2023; Liu et al., 2023). mPLUG-Owl2 (Ye et al., 2023) employs a modality-adaptive language decoder to handle different modalities within distinct modules, which mitigates the issue of modality interference. Given the importance of effectively guiding the model through textual prompts to elicit the desired output, we have selected mPLUG-Owl2 as our feature extractor.
@@
Consequently, we try to propose the rational design of prompts leading LMMs to obtain those image semantic content. mPLUG-Owl2 possess the ability to understand fine-grained semantic contents, but without carefully designed input prompts, some prompts, such as ”Please evaluate if the image quality is compromised due to violations of common human sense or logic?” although expresses the desire for the model to assess whether the semantic content of the image contradicts human perception, would lead to unsatisfactory results. So we meticulously designed two prompts to guide the LMM, denoted as and respctively. Specifically,
@@
corresponding to the existence of semantic content and the coherence of semantic content in images. Test results, as shown in Figure 4 using the mPLUG-Owl2 official demo111https://modelscope.cn/studios/iic/mPLUG-Owl2/summary, have proven these questions to be effective.
However, the text format is not conducive to being utilized by MANIQA to impart semantic insights. To bridge this gap, it’s essential to transform the information into a format that MANIQA can easily leverage. So we extract features from the final layer of mPLUG-Owl2’s hidden layers, achieving an accessible embedded representation of the LMM’s output. This output is a tensor with dimensions of [token_length, hidden_size], where ”token_length” represents the number of output tokens, and ”hidden_size” denotes the dimensionality of the hidden layer representations associated with each token. For mPLUG-Owl2, the hidden_size is set to 4096. Subsequently, we conduct an averaging operation across the token dimension, yielding a vector with dimensions 1x4096. This vector then serves as the basis for further feature fusion procedures. The process can be represented as Equation (3) :
@@ 3.3. Adaptive Fusion Module
Given the complex influence of color, composition, details, semantic content, and other factors on image quality, simply concatenating the extracted features may not always yield the best results. To dynamically fuse a variety of complementary features, we propose the adaptive fusion module (AFM) for organic feature integration. This process can be divided into two main parts. The first part involves transforming the extracted features into a unified vector space of the same dimension, allowing for vector fusion operations. Specifically, for quality-aware features, this transformation block simply applies a Fc layer, transforming them to the same dimension as the original features (1x784). For fine-grained semantic features, it uses a Fc layer to project them onto a 1x784 dimension, followed by a relu activation layer and a dropout layer to enhance the network’s expressive power and generalization. The second part employs a MoE to dynamically fuse the three features. The MoE’s gating network takes the transformed three features as input and outputs dynamic weights , corresponding to three features’ contributions to image quality. Structurally, this gating network comprises a Fc layer and a sigmoid layer. The final image quality representation vector can be obtained through a weighted sum of the three feature vectors. Following the denotation which sign the three features as , , , this process can be represented as:
@@
Type | Method | AGIQA-3k | AIGCQA-20K | KonIQ-10k | SPAQ | ||||
SRCC | PLCC | SRCC | PLCC | SRCC | PLCC | SRCC | PLCC | ||
Handcrafted feature-based | BRISQUE (Mittal et al., 2012a) | 0.4726 | 0.5612 | 0.1663 | 0.3580 | 0.7050 | 0.7070 | 0.8021 | 0.8056 |
NIQE (Mittal et al., 2012b) | 0.5236 | 0.5668 | 0.2085 | 0.3378 | 0.5512 | 0.4881 | 0.7030 | 0.6700 | |
ILNIQE (Zhang et al., 2015) | 0.6097 | 0.6551 | 0.3359 | 0.4551 | 0.4531 | 0.4676 | 0.7194 | 0.6540 | |
LMM-based | CLIPIQA (Wang et al., 2023) | 0.6524 | 0.6968 | 0.4147 | 0.6459 | 0.6957 | 0.7270 | 0.7385 | 0.7354 |
CLIPIQA+ (Wang et al., 2023) | 0.6933 | 0.7493 | 0.4553 | 0.6682 | 0.8756 | 0.8786 | 0.8590 | 0.8796 | |
Q-Align (Wu et al., 2023a) | 0.6728 | 0.6910 | 0.6743 | 0.6815 | 0.9223 | 0.9113 | 0.8875 | 0.8860 | |
Traditional DNN-based | HyperIQA (Su et al., 2020) | 0.8509 | 0.9049 | 0.8162 | 0.8329 | 0.9041 | 0.9152 | 0.9155 | 0.9188 |
MANIQA (Yang et al., 2022) | 0.8618 | 0.9115 | 0.8507 | 0.8870 | 0.9300 | 0.9460 | 0.9229 | 0.9278 | |
DBCNN (Zhang et al., 2018) | 0.8263 | 0.8900 | 0.8054 | 0.8483 | 0.8443 | 0.8624 | 0.9099 | 0.9273 | |
StairIQA (Sun et al., 2023) | 0.8343 | 0.8933 | 0.7899 | 0.8428 | 0.9209 | 0.9362 | 0.9238 | 0.9273 | |
BAID (Yi et al., 2023) | 0.1304 | 0.2030 | 0.1652 | 0.1483 | 0.2535 | 0.2658 | 0.3256 | 0.3232 | |
MUSIQ (Ke et al., 2021) | 0.8261 | 0.8657 | 0.8329 | 0.8646 | 0.8249 | 0.9376 | 0.8732 | 0.8688 | |
DNN with LMM | MA-AGIQA | 0.8939 | 0.9273 | 0.8644 | 0.9050 | 0.9339 | 0.9482 | 0.9278 | 0.9321 |